OpenStudy (anonymous):

0 0 0 0 2 0 1 0 2 3 0 0 0 3 4 1 2 3 4 5 0 0 1 2 3 (A) –6 (B) –3 (C) 6 (D) 0

OpenStudy (jackellyn):

question?

OpenStudy (anonymous):

actually i dont knw its frm which topic :(

OpenStudy (jackellyn):

matrices?

OpenStudy (anonymous):

its written in that way as i copy

OpenStudy (anonymous):

i think so !

OpenStudy (anonymous):

how can u plss guide me on this ?

OpenStudy (anonymous):

If a determinant:\[\text{Det}\left[\left( \begin{array}{ccccc} 0 & 0 & 0 & 0 & 2 \\ 0 & 1 & 0 & 2 & 3 \\ 0 & 0 & 0 & 3 & 4 \\ 1 & 2 & 3 & 4 & 5 \\ 0 & 0 & 1 & 2 & 3 \\ \end{array} \right)\right]=-6 \]

OpenStudy (anonymous):

can u tell me how ?

OpenStudy (anonymous):

http://www.analyzemath.com/Tutorial-System-Equations/determinants.html Your expression was calculated using the Mathematica computer program. This is what computers are for.

OpenStudy (amistre64):

you can swap rows; but each time you do, you need to alternate the sign of the resulting det

OpenStudy (amistre64):

0 0 0 0 2 < 0 1 0 2 3 0 0 0 3 4 1 2 3 4 5 < 0 0 1 2 3 1 2 3 4 5 0 1 0 2 3 0 0 0 3 4 < 0 0 0 0 2 0 0 1 2 3 < 1 2 3 4 5 0 1 0 2 3 0 0 1 2 3 0 0 0 0 2 < 0 0 0 3 4 < 1 2 3 4 5 0 1 0 2 3 0 0 1 2 3 0 0 0 3 4 0 0 0 0 2 that was 3 swaps; -+-, the determinant of this will be the negative of the original

OpenStudy (amistre64):

since we have zeros all under the diagonal; the det is just the product of the diag 1.1.1.3.2 = 6 we want the opposite of 6: -6